OpenStudy (jazzyluvsu123):

PLz Help for a Free MEDAL!!!! Samuel can type nearly 40 words per minute. Use this information to find the number of hours it would take him to type 2.6 x 10 exponent5 scientific Notation! Is my math lesson plz help !

OpenStudy (anonymous):

40 words a minute is \(40\times 60=2400=2.4\times 10^3\) words an hour

OpenStudy (anonymous):

your job it to compute \[\frac{2.6\times 10^5}{2.4\times 10^3}\]

OpenStudy (jazzyluvsu123):

2.4 times 10^3 ??? @satellite73

OpenStudy (anonymous):

that is what 40 words a minute is in words per hour

OpenStudy (jazzyluvsu123):

So I Multiple those Numbers How Would I Change 10^3 Into something Multipliable?

OpenStudy (anonymous):

don't multiply, divide

OpenStudy (jazzyluvsu123):

I Don't Understand

OpenStudy (anonymous):

lets go slow, but then quick

OpenStudy (jazzyluvsu123):

ok

OpenStudy (anonymous):

if you can type 2400 words an hour, then in two hours you can type 4800 words, and in three hours you can type \(3\times 2400\) and in 4 hours you can type \(4\times 2400\) words

OpenStudy (anonymous):

you want to know how long to type \(2.6\times 10^5=260,000\) words

OpenStudy (jazzyluvsu123):

Yes

OpenStudy (anonymous):

to do that you need to divide \[260,000\div 2400\]

OpenStudy (anonymous):

in scientific notation that is \[\frac{2.6\times 10^5}{2.4\times 10^3}\] divide the numbers in front, subtract the exponents

OpenStudy (jazzyluvsu123):